### v1

@@ Introduction
The rapid expansion of social media has fundamentally changed the way people communicate, share information, and consume content. These platforms are now central to society, serving as tools for dissemination, entertainment, and consumption of information [1, 2, 3, 4]. However, their engagement-driven business models and the dynamics of online interactions have raised concerns about their broader social impacts [5, 6], including their role in deepening polarization [7, 8, 9], propagating misinformation [10, 11, 12], and amplifying hate speech [13, 14, 15, 16, 17].
@@
Building on these concerns, today’s digital environment is characterized by an overabundance of information, offering users an unprecedented range of choices for media consumption [18]. This shift results from the proliferation of various sources of content, fueling intense competition for user attention [19, 20, 21], amplified by the engagement-driven strategies of social media platforms. Consequently, the information ecosystem has undergone a significant transformation, leading to a fragmented audience willing to participate [22, 23] but dispersed across numerous platforms and personalized feeds [24]. Traditional media gatekeepers, once central to public discourse, no longer unilaterally set the agenda [25], diverging from previously dominant patterns of centralized media consumption [26, 27, 28, 29]. This radical transformation has led to a vast and unstructured flow of content [30] that may induce information overload [31]. Users often feel overwhelmed by the sheer volume of content when seeking information [32], complicating their ability to navigate this landscape.
@@
In this information-saturated ecosystem, online users increasingly exhibit selective consumption behaviors, engaging primarily with content aligned with their existing beliefs [33, 34, 12] while ignoring opposing viewpoints [35, 36]. This behavior reinforces echo chambers and exacerbates polarization [37, 38, 8], in an environment where the value of news content is no longer determined solely by its intrinsic quality [39]. News consumption has become a shared social experience, with individuals exchanging links and recommendations within their networks and treating news as a form of cultural currency [40]. As a result, traditional hierarchical models of media influence struggle to explain the dynamics of public discourse in this decentralized environment [41, 42, 43], where narratives and participants compete for prominence.
Understanding these processes requires a change in perspective. Classical theories, such as agenda-setting, emphasize the transfer of issue salience from news media to the public agenda [44, 45]. However, these theories fall short of capturing the decentralized, interactive nature of digital ecosystems [46, 27], which have become the new frontier for exploring agenda-setting effects [47]. Concepts like “agendamelding” highlight how users integrate content from multiple sources to construct personalized narratives [48, 49], reflecting the decentralization of influence. Despite the democratizing potential of this bottom-up process, questions remain about what drives public discourse in this fragmented environment.
@@
Using multiple large-scale Twitter (now X) datasets, our analysis outlines the drivers of public discourse by examining three global debates—Climate Change (COP26), COVID-19, and the Russo-Ukrainian War. We compare various actor types, such as media outlets, activists, and individual influencers, to determine whether specific categories dominate discussions and what drives online debates.
@@
Our findings reveal a fragmented environment where no single category of actor consistently dominates discussions. Furthermore, users cluster into two polarized ideological communities, one representing the majority of users endorsing a mainstream narrative and a minority usually embracing alternative views. Notably, a user’s stance in one debate is a reliable marker of their opinion on other topics. For instance, users supporting climate action in the COP26 debate are likely to advocate for vaccination during COVID-19 and aid for Ukraine, while opponents of COVID-19 vaccination are typically climate skeptics and critics of military aid.
@@
These results underscore that online discussions are shaped by shared ideological alignments, rather than by a top-down model dominated by traditional gatekeepers. This highlights how decentralized, bottom-up dynamics dominate public discourse, where ideological structures are pivotal in predicting the evolution of debates and the spread of specific narratives. By identifying consistent patterns in user engagement and ideological alignment, this study provides insights into the mechanisms shaping public discourse in an information-saturated landscape.
@@
Each influencer is then manually assigned to one of six categories: Activist, International Organization/NGO, Media, Politics, Private Individual, and Other. Accounts whose tweets are unrelated to the topics of discussion are excluded from the analysis. In addition to categorizing these accounts according to their role, we annotate each account with their stance on key issues within the debates. The final sets of influencers are comprised of 561 accounts for COP26, 1577 for COVID-19, and 1166 for Ukraine. The labeling process is described more in detail in Methods and in Supplementary Section 1.1.
To understand whether the engagement that influencers receive depends on the category they belong to, we analyze the distribution of retweets received by each category per debate. This analysis enables us to assess whether certain categories of influencers are more effective in driving users’ engagement.
@@
To statistically test these insights, we perform pairwise comparisons between the categories within each debate using Kolmogorov-Smirnov tests, which assess whether the retweet distributions for any two categories are drawn from the same population. The results, detailed in Supplementary Section 1.3, show no statistically significant differences between distributions, suggesting that the category of an influencer does not affect the number of retweets they receive. Additional results, including the prevalence of each influencer category across retweet count buckets, are provided in Supplementary Information.
@@
We compare the observed and randomized distributions using two-sided two-sample Wilcoxon rank-sum tests for both strategies. The results show no significant differences, indicating that the observed number of retweets is consistent with random assignment of categories. This reinforces the finding that an actor’s category does not appear to affect their ability to attract retweets. Detailed results, including visualizations of the randomized distributions, are provided in Supplementary Section 1.3.
Finally, we analyze users’ preferences in retweeting influencer categories. To ensure that all users in a dataset might have retweeted each category at least once, we restrict our analysis to users with at least six retweets and compute how frequently they retweet each category. We then employ the Gini Index (GI)  [50] to measure how evenly users distribute their retweets across influencer categories. The GI is a value in $[0,1]$ that quantifies inequality within a distribution: a value of 00 corresponds to minimum inequality (i.e., all categories are retweeted by the user with the same frequency), while a value of $1$ corresponds to maximum inequality (i.e., the user only retweets a single category). See Methods for its formal definition.
The resulting distributions of users’ GI, displayed in Supplementary Fig. S5, suggest that users show a preference for a limited subset of categories. For COP26, the GI distribution has a median of $0.73$ (interquartile range: $Q1=0.60,\ Q3=0.87$); for COVID-19, the median is $0.77$ ($Q1=0.65,\ Q3=0.89$); and for Ukraine, the median is $0.69$ ($Q1=0.58,\ Q3=0.80$). However, when considering the total number of retweets each category receives, we observe lower levels of inequality across all debates: $\text{GI}=0.28$ for COP26, $\text{GI}=0.59$ for COVID-19, and $\text{GI}=0.53$ for Ukraine.
@@
We begin by constructing an interaction matrix where each element corresponds to the number of times a user retweets a specific influencer. Then, the method assigns numerical scores to users so that individuals retweeting similar subsets of influencers receive similar scores. By rescaling these values to the range $[-1,+1]$, we obtain an ideological spectrum, with users at opposing ends interacting with distinct and non-overlapping sets of influencers. A detailed explanation of the technique is provided in Methods.
Following the methodology adopted in previous works [51], we focus on the top 300 most retweeted influencers from each of the three debates. We calculate the latent ideology of all users within each debate and validate the resulting scores to ensure that the distributions represent true ideological spectra (see Supplementary Section 2.1). Finally, we follow standard procedure and estimate the ideological position of each influencer as the median ideology score of their retweeters [51].
Figure 2 illustrates the latent ideology distributions for users and influencers across the three debates. To ensure consistency, we map the majority group to $-1$ and the minority group to $+1$ for all cases. Although the extent of polarization varies between debates, all distributions exhibit a clear bimodal structure, with most users clustered near the spectrum’s extremes. This observation indicates the existence of two dominant, ideologically opposing communities within each debate.
@@
Figure 2: Ideological spectra of users and influencers in the COP26, COVID-19, and Ukraine debates. The majority group is mapped to $-1$, whereas the minority group to $+1$. All distributions exhibit bimodality, indicating the presence in each debate of two main communities of users who retweet distinct sets of influencers.
@@
The results, presented in Fig. 4, further support the qualitative insights derived from the joint distribution in Fig. 3. The two matrices display the fraction of users who remain in the majority (or minority) group in a debate on the $x$-axis, given that they belong to the majority (or minority) group in the debate specified on the $y$-axis. This asymmetry arises from the different fractions of majority and minority supporters in each debate. For instance, the proportion of COP26 supporters who also endorse COVID-19 vaccines may differ from the proportion of vaccine supporters who also back climate change policies. The matrices reveal that a user’s stance in one debate is highly indicative of their stance in others, with the proportion of users maintaining the same ideological leaning equaling or exceeding 90% for most debate pairs. This result is robust to using more restrictive ideology thresholds to label users as ‘Majority’ or ‘Minority’ (see Supplementary Section 2.2).
@@ Conclusions
Our study provides new insights into the interplay between user engagement and influencers’ role in major global debates on social media. By analyzing large-scale datasets from the COP26, COVID-19, and Russo-Ukrainian War debates, we demonstrate that user ideological alignment holds greater significance than an influencer’s category in shaping online discussions.
In traditional media environments, agenda-setting theories emphasize the hierarchical influence of institutional actors, such as mainstream media and political elites, in prioritizing public issues. Our findings suggest that these dynamics are significantly transformed in the context of social media. While certain users achieve prominence within specific topics, their relevance does not appear to be tied to their social or professional role, nor is it consistent across debates. Rather, the engagement they receive may arise from resonance within polarized communities.
These findings point out a shift from the top-down agenda-setting model to decentralized, bottom-up mechanisms, where users collectively amplify narratives aligned with their ideological preferences. The persistence of a Majority versus Minority ideological configuration in all debates highlights the presence of segregated and ideologically contrasting communities that extend beyond specific topics. This consistent affiliation may further limit exposure to diverse perspectives, thereby reinforcing the effects of echo chambers. Further, despite the difference in the size of their communities, the comparable engagement levels obtained by Majority and Minority influencers suggest that their popularity is not directly related to the ideological stance they assume.
@@
The decentralized nature of social media discourse fostered a new model of information spreading, challenging the traditional gatekeeping power of dominant actors. This new paradigm presents both opportunities and risks. On one hand, it democratizes the flow of information, amplifying diverse perspectives. On the other hand, it risks exacerbating polarization and misinformation, as users selectively engage with ideologically congruent content across topics.
Despite the robustness of our analyses, several limitations must be acknowledged. First, our investigation focuses on Twitter and may not fully capture user dynamics on other platforms. Additionally, the classification of users into Majority and Minority groups relies on latent ideology modeling, which, while effective, may simplify the multifaceted nature of user beliefs. Future research should extend these analyses to explore cross-platform dynamics, temporal shifts in ideological alignment, and the role of emerging fringe platforms in shaping discourse.

### v2

@@ Introduction
The rapid expansion of social media has fundamentally changed the way people communicate, share information, and consume content. These platforms are now central to society, serving as tools for dissemination, entertainment, and consumption of information [1, 2, 3, 4]. However, their engagement-driven business models and the dynamics of online interactions have raised concerns about their broader social impacts [5, 6], including their role in deepening polarization [7, 8, 9], propagating misinformation [10, 11, 12], and amplifying hate speech [13, 14, 15, 16, 17]. Moreover, recent research has highlighted the central role that the structure of social networks has in disseminating information [18].
@@
Building on these concerns, today’s digital environment is characterized by an overabundance of information, offering users an unprecedented range of choices for media consumption [19]. This shift results from the proliferation of various sources of content, fueling intense competition for user attention [20, 21, 22], amplified by the engagement-driven strategies of social media platforms. Consequently, the information ecosystem has undergone a significant transformation, leading to a fragmented audience willing to participate [23, 24] but dispersed across numerous platforms and personalized feeds [25]. Traditional media gatekeepers, once central to public discourse, no longer unilaterally set the agenda [26], diverging from previously dominant patterns of centralized media consumption [27, 28, 29, 30].
@@
This radical transformation has led to a vast and unstructured flow of content [31] that may induce information overload [32]. Users often feel overwhelmed by the sheer volume of content when seeking information [33], complicating their ability to navigate this landscape. In this information-saturated ecosystem, online users increasingly exhibit selective consumption behaviors, engaging primarily with content aligned with their existing beliefs [34, 35, 12] while ignoring opposing viewpoints [36, 37]. This behavior reinforces echo chambers and exacerbates polarization [38, 39, 8], in an environment where the value of news content is no longer determined solely by its intrinsic quality [40]. News consumption has become a shared social experience, with individuals exchanging links and recommendations within their networks and treating news as a form of cultural currency [41]. As a result, traditional hierarchical models of media influence struggle to explain the dynamics of public discourse in this decentralized environment [42, 43, 44], where narratives and participants compete for prominence.
Previous studies have explored how social media have changed the traditional gatekeeping function of legacy media. Contemporary debates on social platforms are frequently shaped not by institutional actors, but by networked publics [45] participating in the online public sphere. In classic media research, gatekeeping was understood as the process through which mass media filtered, shaped, and prioritized information flows to audiences [46, 47]. However, the rise of participatory digital technologies has profoundly challenged this top-down model. Scholars have argued that social media has empowered users to actively contribute to the creation and dissemination of news content, thereby eroding the media’s monopoly on gatekeeping [48, 49]. As Bro and Wallberg[50] note, individuals increasingly access news through peer networks, such as family and friends, rather than through editorially curated channels. This shift has altered not only how news spreads, but also who holds the power to define it. While news organizations continue to seek authority over informational flows [51], alternative actors have emerged as influential newsmakers in the digital sphere. This transformation has led some scholars to reconceptualize gatekeeping as a distributed, collaborative process within communicative networks. Hoskins and O’Loughlin [52] initially questioned whether gatekeeping could be meaningfully applied to such networked contexts, but their research—along with subsequent studies—supports the notion of “network gatekeeping,” a dynamic process emerging from diffused and collaborative gatekeeping processes [52, 53]. Our findings contribute to this literature by illustrating how agenda-setting processes might be increasingly negotiated among a wide array of actors, in an environment where news also emerges from bottom-up dynamics.
While classical agenda-setting theory, originating with McCombs and Shaw [54], was built around a top-down logic typical of mass communication, subsequent scholarship has explored its evolution in light of media digitalization, debating the relevance of the theory in digital settings. The same authors, on the 50th anniversary of the theory, traced its evolution by identifying seven distinct facets, emphasizing need for orientation, network agenda setting, and agendamelding as particularly active areas of contemporary research [55, 56] within an increasingly complex and interactive media ecosystem [57]. Some scholars have argued that agenda setting effects remain operative, albeit adapted to the new media landscape [58], while others have pointed to a weakening of traditional media’s influence in favor of digital platforms [59, 60], or have conceptualized social media themselves as a further development of agenda setting mechanisms [61]. Russell Neuman et al. [62] identified several patterns of issue framing in social and traditional media that highlighted an interdependence of the two rather than a one-way agenda-setting dynamic. Feezell [63], instead, investigated how social media contributes to the dissemination of mainstream news and their perceived importance. New studies on agenda setting processes have also kept in consideration various platforms, such as Twitter [64, 65, 66], Instagram [67] and Reddit [68]. Analyzing 71.77 million tweets, Yi and Wang [69] found that professional media influence only a small number of individuals, while they argue that opinion leaders have become strong competitors of traditional professional media in agenda setting processes.
@@
Understanding these processes requires a change in perspective. Using multiple large-scale Twitter (now X) datasets, our analysis outlines the drivers of public discourse by examining three global debates—climate change (COP26), COVID-19, and the Russo-Ukrainian War. We compare various actor types, such as media outlets, activists, and individual influencers, to determine whether specific categories dominate discussions and what drives online debates. Furthermore, our diverse selection of debates, spanning environmental, health, and geopolitical domains, enables us to assess whether the dynamics we investigate are tied to specific topics or reflect systemic patterns across issues.
@@
In particular, we aim to address the following questions: What are the drivers of user engagement in fragmented social media environments? Do actor categories or ideological alignment better explain how attention is allocated? Our findings reveal a fragmented environment where no single category of actor consistently dominates discussions. Furthermore, users cluster into two polarized ideological communities, one representing the majority of users endorsing a mainstream narrative and a minority usually embracing alternative views. Notably, a user’s stance in one debate is a reliable marker of their opinion on other topics. For instance, users supporting climate action in the COP26 debate are likely to advocate for vaccination during COVID-19 and aid for Ukraine, while opponents of COVID-19 vaccination are typically climate skeptics and critics of military aid. This consistency in ideological stance across heterogeneous debates suggests that polarization on social media emerges from overarching ideological divides, rather than being confined to specific issues.
@@
Our results underscore that online discussions are shaped by shared ideological alignments, rather than by a top-down model dominated by traditional gatekeepers. This highlights how decentralized, bottom-up dynamics dominate public discourse, where ideological structures are pivotal in predicting the evolution of debates and the spread of specific narratives. By identifying consistent patterns in user engagement and ideological alignment, this study provides insights into the mechanisms shaping public discourse in an information-saturated landscape.
@@
Each influencer is then manually assigned to one of six categories: Activist, International Organization/NGO, Media, Politics, Private Individual, and Other. Accounts whose tweets are unrelated to the topics of discussion are excluded from the analysis. In addition to categorizing these accounts according to their role, we annotate each account with their stance on key issues within the debates. The final sets of influencers are comprised of 561 accounts for COP26, 1577 for COVID-19, and 1166 for Ukraine. The labeling process is described more in detail in Methods and in Supplementary Information, where we also report the proportion of influencers in each category (Supplementary Fig. S1) and the distributions of original tweets posted by these accounts (Supplementary Fig. S2).
One fundamental question is the following: according to agenda-setting theory, we might expect media or political accounts to receive higher levels of engagement than other types of information providers. To understand if this is the case, namely whether the engagement that influencers receive depends on the category they belong to, we analyze the distribution of retweets received by each category per debate. This analysis enables us to assess whether certain categories of influencers are more effective in driving users’ engagement.
@@
To statistically test these insights, we perform pairwise comparisons between the categories within each debate using Kolmogorov-Smirnov tests, which assess whether the retweet distributions for any two categories are drawn from the same population. The results, reported in Supplementary Table S1, show no statistically significant differences between distributions, suggesting that the category of an influencer does not affect the number of retweets they receive. Additional results are provided in Supplementary Information, including the prevalence of each influencer category across retweet count buckets (Supplementary Fig. S3).
@@
We compare the observed and randomized distributions using two-sided two-sample Wilcoxon rank-sum tests for both strategies. The results show no significant differences, indicating that the observed number of retweets is consistent with random assignment of categories. This reinforces the finding that an actor’s category does not appear to affect their ability to attract retweets. Detailed results are provided in Supplementary Table S2, as well as visualizations of the randomized distributions (Supplementary Fig. S4).
Finally, we analyze users’ preferences in retweeting influencer categories. To ensure that all users in a dataset might have retweeted each category at least once, we restrict our analysis to users with at least six retweets and compute how frequently they retweet each category. We then employ the Gini Index (GI) [70] to measure how evenly users distribute their retweets across influencer categories. The GI is a value in $[0,1]$ that quantifies inequality within a distribution: a value of 0 corresponds to minimum inequality (i.e., all categories are retweeted by the user with the same frequency), while a value of $1$ corresponds to maximum inequality (i.e., the user only retweets a single category). See Methods for its formal definition.
The resulting distributions of users’ GI, displayed in Supplementary Fig. S5, suggest that users show a preference for a limited subset of categories. For COP26, the GI distribution has a median of $0.73$ (interquartile range: Q1=0.60,Q3=0.87Q1=0.60,\ Q3=0.87); for COVID-19, the median is $0.77$ (Q1=0.65,Q3=0.89Q1=0.65,\ Q3=0.89); and for Ukraine, the median is $0.69$ (Q1=0.58,Q3=0.80Q1=0.58,\ Q3=0.80). However, when considering the total number of retweets each category receives, we observe lower levels of inequality across all debates: $\text{GI}=0.28$ for COP26, $\text{GI}=0.59$ for COVID-19, and $\text{GI}=0.53$ for Ukraine.
@@
We begin by constructing an interaction matrix where each element corresponds to the number of times a user retweets a specific influencer. Then, the method assigns numerical scores to users so that individuals retweeting similar subsets of influencers receive similar scores. By rescaling these values to the range [1,+1][-1,+1], we obtain an ideological spectrum, with users at opposing ends interacting with distinct and non-overlapping sets of influencers. A detailed explanation of the technique is provided in Methods.
Following the methodology adopted in previous works [71], we focus on the top 300 most retweeted influencers from each of the three debates. We calculate the latent ideology of all users within each debate and validate the resulting scores to ensure that the distributions represent true ideological spectra (see Supplementary Section 2.1 and Supplementary Fig. S6). Finally, we follow standard procedure and estimate the ideological position of each influencer as the median ideology score of their retweeters [71].
Figure 2 illustrates the latent ideology distributions for users and influencers across the three debates. To ensure consistency, we map the majority group to 1-1 and the minority group to +1+1 for all cases. Although the extent of polarization varies between debates, all distributions exhibit a clear bimodal structure, with most users clustered near the spectrum’s extremes. This observation indicates the existence of two dominant, ideologically opposing communities within each debate.
@@
Figure 2: Ideological spectra of users and influencers in the COP26, COVID-19, and Ukraine debates. The majority group is mapped to 1-1, whereas the minority group to +1+1. All distributions exhibit bimodality, indicating the presence in each debate of two main communities of users who retweet distinct sets of influencers.
@@
The results, presented in Fig. 4, further support the qualitative insights derived from the joint distribution in Fig. 3. The two matrices display the fraction of users who remain in the majority (or minority) group in a debate on the $x$-axis, given that they belong to the majority (or minority) group in the debate specified on the $y$-axis. This asymmetry arises from the different fractions of majority and minority supporters in each debate. For instance, the proportion of COP26 supporters who also endorse COVID-19 vaccines may differ from the proportion of vaccine supporters who also back climate change policies. The matrices reveal that a user’s stance in one debate is highly indicative of their stance in others, with the proportion of users maintaining the same ideological leaning equaling or exceeding 90% for most debate pairs. This result is robust to using more restrictive ideology thresholds to label users as ‘Majority’ or ‘Minority’ (Supplementary Fig. S8).
@@ Conclusions
Our study provides new insights into the interplay between user engagement and influencers’ roles in major global debates on social media. By analyzing large-scale datasets from the COP26, COVID-19, and Russo-Ukrainian War debates, we demonstrate that user ideological alignment holds greater significance than an influencer’s category in shaping online discussions.
In traditional media environments, agenda-setting theories emphasize the hierarchical influence of institutional actors, such as mainstream media and political elites, in prioritizing public issues. Our findings suggest that these dynamics may not fully extend to social media. While certain users achieve prominence within specific topics, their relevance does not appear to be tied to their social or professional role, nor is it consistent across debates. Rather, the engagement they receive may arise from resonance within polarized communities.
These findings hint at a change in the traditional top-down agenda setting model, pointing to a growing role for decentralized, bottom-up mechanisms, where users collectively amplify narratives aligned with their ideological preferences. The persistence of a Majority versus Minority ideological configuration in all debates highlights the presence of segregated and ideologically contrasting communities that extend beyond specific topics. This consistent affiliation may further limit exposure to diverse perspectives, thereby reinforcing the effects of echo chambers. Further, despite the difference in the size of their communities, the comparable engagement levels obtained by Majority and Minority influencers suggest that their popularity is not directly related to the ideological stance they assume.
@@
Our results also offer some noteworthy applications. For communication professionals, they suggest a need to rethink outreach strategies: rather than targeting audiences solely based on topic relevance or demographic features, it may be more effective to account for users’ broader ideological positioning. Tailoring messages that bridge ideological narratives or are endorsed by trusted influencers within specific communities could increase the chance of cross-cluster diffusion. However, the susceptibility of this type of dynamic shall be taken into account also from a policymaking point of view, as it may fuel the cross-spreading of radicalized narratives between different sensitive topics. Additionally, our findings contribute to the understanding of filter bubbles by demonstrating that they are not isolated to specific issues but reflect systemic, cross-topic ideological alignment. This presents challenges for initiatives aimed at promoting exposure to diverse perspectives. Efforts to mitigate filter bubbles might benefit from designing interventions that operate across topics and target underlying ideological structures, rather than focusing narrowly on specific debates.
The decentralized nature of social media discourse fostered a new model of information spreading, challenging the traditional gatekeeping power of dominant actors. This new paradigm presents both opportunities and risks. On one hand, it democratizes the flow of information, amplifying diverse perspectives. On the other hand, it risks exacerbating polarization and misinformation, as users selectively engage with ideologically congruent content across topics. Moreover, as the traditional filters that once ensured a baseline of editorial standards have weakened, questions arise about the quality of content consumed in online environments. In turn, the fragmentation of the information space undermines the conditions for a shared public debate, an essential foundation of any democratic society.
Despite the robustness of our analyses, several limitations must be acknowledged. First, our investigation focuses exclusively on Twitter and may not fully reflect user dynamics on other platforms, as platform-specific biases and algorithmic curation could, in principle, influence the patterns we observe. Additionally, the classification of users into Majority and Minority groups relies on latent ideology modeling, which, while effective, may simplify the multifaceted nature of user beliefs.
At the same time, we only focused on English tweets, which has to be addressed as a constraint on the generalizability of the findings. However, this detail also underscores a solidity in our findings. Despite focusing on a language used predominantly by non-local audiences, we still observe strong patterns of ideological alignment and cross-topic coherence. For instance, regarding the Russo-Ukrainian debate, even within English-speaking communities—often external to the conflict zone—polarized narratives persist and align with ideological stances in other unrelated debates (e.g., climate change and COVID-19). This upshot suggests that the ideological dynamics we identify are not only local or topic-specific but reflect a broader structure of online polarization that transcends geographic and linguistic boundaries. Nevertheless, we acknowledge that expanding future analyses to include multilingual datasets could capture more nuances in local perspectives and offer a more comprehensive view of discourse dynamics within involved regions.
Future research should extend these analyses to explore cross-platform dynamics, temporal shifts in ideological alignment, and the role of emerging fringe platforms in shaping discourse.